OpenStudy (anonymous):

what is the lenght of a area of 675?

OpenStudy (anonymous):

btw it is a square

OpenStudy (raden):

a^2 = 675 solve for a by taking the square root of 675 then simplify

OpenStudy (raden):

a^2 = 675 a = sqrt(675) = sqrt(225 * 3) = sqrt(225) * sqrt(3) = 15 sqrt(3)